# Instantaneous Type I blow-up and non-uniqueness of smooth solutions of the Navier–Stokes equations

### v1

@@ 4.1. Error bounds on the velocity
$F_{j,j^{\prime},k}$ $=\Delta(\psi_{j,k}-N_{j,k}^{-2}a_{j,k}\varphi_{j,k}\theta_{j}\sin(N_{j,k}\eta_{j}\cdot x))\otimes\Delta\psi_{j^{\prime},k}$
$+$ $\Delta(N_{j,k}^{-2}a_{j,k}\varphi_{j,k}\theta_{j}\sin(N_{j,k}\eta_{j}\cdot x))\otimes\Delta(\psi_{j^{\prime},k}-N_{j^{\prime},k}^{-2}a_{j^{\prime},k}\varphi_{j^{\prime},k}\theta_{j^{\prime}}\sin(N_{j^{\prime},k}\eta_{j^{\prime}}\cdot x)),$
and use the fact that $\|\nabla^{n}(f-\phi_{k}*f)\|_{\infty}\lesssim\ell_{k}\|\nabla^{n+1}f\|_{\infty}$ along with (3.11) to obtain
$\|\nabla^{n}\Delta(\psi_{j,k}-N_{j,k}^{-2}a_{j,k}\varphi_{j,k}\theta_{j}\sin(N_{j,k}\eta_{j}\cdot x))\|_{\infty}$
$\qquad\lesssim N_{j,k}^{-2}\ell_{k}\|\nabla^{3+n}(a_{j,k}\varphi_{j,k}\sin(N_{j,k}\eta_{j}\cdot x))\|_{\infty}$
$\qquad\lesssim N_{j,k}^{1+n}\ell_{k}.$
@@
$\|\nabla^{n}\Delta\psi_{j,k}\|_{\infty}+\|\nabla^{n}\Delta(N_{j,k}^{-2}a_{j,k}\varphi_{j,k}\theta_{j}\sin(N_{j,k}\eta_{j}\cdot x))\|_{\infty}$ $\lesssim N_{j,k}^{n}.$
@@
$\|\nabla^{n}\mathbb{Q}F_{k,j,j^{\prime}}\|_{\infty}$ $\lesssim N_{j\vee j^{\prime},k}^{n+1+\varepsilon}\ell_{k},$
@@
$\|\nabla^{n}\mathcal{E}_{k}^{3}\|_{\infty}$ $\lesssim\sum_{j,j^{\prime}}N_{j,k+1}N_{j^{\prime},k+1}N_{j\vee j^{\prime},k+1}^{-2}N_{j\vee j^{\prime},k}^{n+1+\varepsilon}\ell_{k}\lesssim\sum_{j}N_{j,k}^{n+1+\varepsilon}\ell_{k}.$
@@
$\begin{split}\|\nabla^{n}(\mathcal{J}_{k}^{1}+\mathcal{J}_{k}^{2}+\mathcal{E}_{k})\|_{\infty}&\leq\|\nabla^{n}\mathcal{J}_{k}^{1}\|_{\infty}+\|\nabla^{n}\mathcal{J}_{k}^{2}\|_{\infty}+\|\nabla^{n}\mathcal{E}_{k}^{1}\|_{\infty}+\|\nabla^{n}\mathcal{E}_{k}^{2}\|_{L^{\infty}}+\|\nabla^{n}\mathcal{E}_{k}^{3}\|_{\infty}\\ &\lesssim_{\varepsilon,m}\sum_{j}N_{j,k+1}^{n-1+\varepsilon}N_{J_{d},k}\big(e^{-\frac{1}{4}M_{j,k+1}^{2}t}+M_{j,k+1}^{-m}N_{J_{d},k}^{m}\big)\\ &\quad\,\,+\sum_{j^{\prime}<j}N^{n-1+\varepsilon}_{j,k+1}N_{j^{\prime},k+1}e^{-\frac{1}{4}N_{j,k+1}^{2}t}\\ &\quad\,\,+\sum_{j^{\prime}<j}N_{j,k+1}^{-1}N_{j^{\prime},k+1}N_{j,k+1}^{n-m}M_{j,k+1}^{m+\varepsilon}\\ &\quad\,\,+\sum_{j}\big(N_{j,k+1}^{n-1+\varepsilon}e^{-\frac{1}{4}N_{j,k+1}^{2}t}+N_{j,k+1}^{-2}M_{j,k+1}^{n+2+\varepsilon}\big)+\sum_{j}N_{j,k}^{n+1+\varepsilon}\ell_{k}.\end{split}$
@@
for $b$ and $m$ large enough. For $n=0$ the bound becomes worse on $0,t_{k}]$, as in Proposition 4.1. ∎
@@ 4.2. Bounds on the solution and residual
for some $\alpha,\kappa\in(0,1)$. Moreover,
@@
$\lesssim t^{-1-\frac{n}{2}+\frac{1}{2}(1-b^{1/J_{d}}-4\kappa)}.$
@@
$J_{2}=$ $\sum_{j}\Big(\sum_{k:N_{j,k}<t^{-1/2}}\int_{t^{\prime}}^{t}N_{j,k}s^{-\frac{1}{2}}ds+\sum_{(t^{\prime})^{-1/2}\leq N_{j,k}\leq t^{-1/2}}\int_{t^{\prime}}^{t}N_{j,k}s^{-\frac{1}{2}}e^{-N_{j,k}^{2}s}ds$
@@
$J_{2}\lesssim 1+\log_{A^{b-1}}(t^{-\frac{1}{2}}/(t^{\prime})^{-\frac{1}{2}})$
@@ 5. Construction of the corrector
Let $U$ be any classical solution to the Navier–Stokes equations on the time interval $0,T_{*}]$ which, once again, we replace with $T_{*},T-T_{*}]$ by translation. We may define
@@
The family of solutions $u^{(\sigma)}$ claimed in Theorem 1.6 will arise by modulating a frequency scale $N_{0}>0$ (see Section 6). Having constructed $v$ on $\mathbb{T}^{d}\times[0,\infty)$, we extend it by 0 to the full time interval $\mathbb{R}$, then rescale it to $v^{N_{0}}$, a $2\pi/N_{0}$-periodic vector field $V$ on $\mathbb{R}^{d}\times[0,\infty)$. Here, for an arbitrary vector field on $\mathbb{R}^{d}$, we define the natural Navier–Stokes rescaling
@@
$\partial_{t}w-\Delta w+\mathbb{P}\operatorname{div}(w\otimes w+2(U^{1/N_{0}}+v)\odot w)=-\mathbb{P}\operatorname{div}f,\quad w|_{t=0}=0.$
@@ 5.1. Semigroup estimates
Then we have the following product rule, which is immediate from the definitions and the estimate $\|fg\|_{C^{1,\kappa}}\lesssim\|f\|_{C^{1,\kappa}}\|g\|_{L^{\infty}}+\|f\|_{L^{\infty}}\|g\|_{C^{1,\kappa}}$.
@@
If $f,g\in X$, then $f\odot g\in Y$ with
$\|f\odot g\|_{Y}\lesssim\|f\|_{X}\|g\|_{X}.$
@@
$\|S(t,t^{\prime})\phi\|_{L^{\infty}(\mathbb{T}^{d})}+(t-t^{\prime})^{\frac{1}{2}}\|\nabla S(t,t^{\prime})\phi\|_{C^{\kappa}(\mathbb{T}^{d})}\lesssim_{\overline{C}}t^{-\frac{1}{2}}(t^{\prime})^{-1+\alpha}(t/t^{\prime})^{\epsilon}\|\phi\|_{Y}$
@@
$\begin{split}S(t,t^{\prime})\phi&=e^{(t-t^{\prime})\Delta}\mathbb{P}\operatorname{div}\phi(t^{\prime})-2\int_{t^{\prime}}^{t}e^{(t-s)\Delta}\mathbb{P}\operatorname{div}(((U^{1/N_{0}}+v)(s)\odot S(s,t^{\prime})\phi)\,ds\\ &=e^{(t-t^{\prime})\Delta}\mathbb{P}\operatorname{div}\phi(t^{\prime})-2\int_{t^{\prime}}^{t^{\prime}\vee{\frac{t}{2}}}e^{(t-s)\Delta}\mathbb{P}\operatorname{div}((U^{1/N_{0}}+v)(s)\odot S(s,t^{\prime})\phi)\,ds\\ &\quad-2\int_{t^{\prime}\vee{\frac{t}{2}}}^{t}e^{(t-s)\Delta}\mathbb{P}\operatorname{div}((U^{1/N_{0}}+v)(s)\odot S(t,t^{\prime})\phi)\,ds\\ &=:I_{1}+I_{2}+I_{3}.\end{split}$
@@
Therefore we need estimate for $\|v\|_{L^{1}([t^{\prime},t],t^{-\frac{1}{2}}dt;L^{\infty})}$ and $\|v\|_{L^{2}([t^{\prime},t];L^{\infty})}$. Using (4.19), we obtain
@@
$\begin{split}\nabla S(t,t^{\prime})\phi&=e^{(t-t^{\prime})\Delta}\nabla\mathbb{P}\operatorname{div}\phi(t^{\prime})-2\int_{t^{\prime}}^{t}e^{(t-s)\Delta}\nabla\mathbb{P}\operatorname{div}((U+v(s))\odot S(t,t^{\prime})\phi)\,ds\\ &=e^{(t-t^{\prime})\Delta}\nabla\mathbb{P}\operatorname{div}\phi(t^{\prime})-2\int_{t^{\prime}}^{\frac{1}{2}(t+t^{\prime})}e^{(t-s)\Delta}\nabla\mathbb{P}\operatorname{div}((U+v(s))\odot S(t,t^{\prime})\phi)\,ds\\ &\quad-2\int_{\frac{1}{2}(t+t^{\prime})}^{t}e^{(t-s)\Delta}\nabla\mathbb{P}\operatorname{div}((U+v(s))\odot S(t,t^{\prime})\phi)\,ds\\ &=:I_{4}+I_{5}+I_{6}.\end{split}$
@@
$\|e^{t\Delta}\nabla^{m}\mathbb{P}f\|_{C^{r}}\lesssim t^{-\frac{m+r-s}{2}}\|f\|_{C^{s}}$ (5.4)
@@ 5.2. Fixed point argument
There exists $\varepsilon>0$ such that for all $\delta\in(0,\varepsilon)$ and $A\gg 1$, there exists $w\in B_{X}(0,\delta)$ with $u^{1/N_{0}}=U^{1/N_{0}}+v+w$ solving (1.1) on $0,\bar{T}]\times\mathbb{R}^{d}$. Further, choosing $\overline{C}$ and $N_{0}$ large depending on $U$ and $T$, the solution $u^{1/N_{0}}$ extends to a classical solution on $[0,T]$.
@@
$F(w)=-\int_{0}^{t}S(t,t^{\prime})(w\otimes w+f(t^{\prime}))dt^{\prime}$
@@
$\|F(w)\|_{X}$ $\leq\sup_{t\in(0,1]}\int_{0}^{t}\Big(t^{\frac{1}{2}(1-\alpha)}\|S(t,t^{\prime})(w\otimes w+f(t^{\prime}))\|_{\infty}$
$\qquad+t^{1-\frac{\alpha}{2}}\|\nabla S(t,t^{\prime})(w\otimes w+f(t^{\prime}))\|_{C^{\kappa}}\Big)dt^{\prime}$
$\lesssim_{\bar{C}}\sup_{t\in(0,1]}\int_{0}^{t}\Big(t^{\frac{1}{2}(1-\alpha)}+t^{1-\frac{\alpha}{2}}(t-t^{\prime})^{-\frac{1}{2}}\Big)t^{-\frac{1}{2}}(t^{\prime})^{-1+\alpha}\left(\frac{t}{t^{\prime}}\right)^{\epsilon}\|w\otimes w+f\|_{Y}dt^{\prime}$
$\lesssim\|w\otimes w\|_{Y}+\|f\|_{Y},$
choosing $\epsilon\in(0,\alpha)$. We conclude by Proposition 4.3 and Lemma 5.1 that
$\|F(w)\|_{X}$ $\leq C(\|w\|_{X}^{2}+\|f\|_{Y})\leq C^{\prime}(\delta^{2}+\epsilon_{0}),$
@@
Taking $\delta>0$ small enough, we conclude the existence of a fixed point $w\in B_{X}(0,\delta)$ of $F$. This proves the claim about $u$ on $0,\bar{T}]$. For $N_{0}$ sufficiently large, it will be the case that $N_{0}^{2}(T-T_{*})>\bar{T}=\bar{C}$. To extend the lifespan, we argue that global solution theory can be applied starting from initial data $u(T_{*})$. Observe that by (4.18) and the fact that $w\in X$, the perturbation $v(\bar{T})+w(\bar{T})$ can be made to be arbitrarily small with the choice of $\bar{T}$. Moreover, $U^{1/N_{0}}(\bar{T})$ stays bounded in $BMO^{-1}$ by the scaling invariance. Taking $\bar{T}$ as large as needed depending on $U$, we conclude global existence of $u$ by (for instance) the stability theorem of Auscher, Dubois, and Tchamitchian [MR2062638]. ∎

### v2

@@ 5.1. Error bounds on the velocity
$F_{j,j^{\prime},k}$ $=\Delta(\psi_{j,k+1}-N_{j,k+1}^{-2}a_{j,k+1}\varphi_{j,k+1}\theta_{j}\sin(N_{j,k+1}\eta_{j}\cdot x))\otimes\Delta\psi_{j^{\prime},k+1}$
$\quad+\Delta(N_{j,k+1}^{-2}a_{j,k+1}\varphi_{j,k+1}\theta_{j}\sin(N_{j,k+1}\eta_{j}\cdot x))$
$\qquad\otimes\Delta(\psi_{j^{\prime},k+1}-N_{j^{\prime},k+1}^{-2}a_{j^{\prime},k+1}\varphi_{j^{\prime},k+1}\theta_{j^{\prime}}\sin(N_{j^{\prime},k+1}\eta_{j^{\prime}}\cdot x)),$
and use the fact that $\|\nabla^{n}(f-\phi_{k+1}*f)\|_{\infty}\lesssim\ell_{k+1}\|\nabla^{n+1}f\|_{\infty}$ along with (4.13) to obtain
$\|\nabla^{n}\Delta(\psi_{j,k+1}-N_{j,k+1}^{-2}a_{j,k+1}\varphi_{j,k+1}\theta_{j}\sin(N_{j,k+1}\eta_{j}\cdot x))\|_{\infty}$
$\qquad\lesssim N_{j,k+1}^{-2}\ell_{k+1}\|\nabla^{3+n}(a_{j,k+1}\varphi_{j,k+1}\sin(N_{j,k+1}\eta_{j}\cdot x))\|_{\infty}$
$\qquad\lesssim N_{j,k+1}^{1+n}\ell_{k+1}.$
@@
$\|\nabla^{n}\Delta\psi_{j,k+1}\|_{\infty}+\|\nabla^{n}\Delta(N_{j,k+1}^{-2}a_{j,k+1}\varphi_{j,k+1}\theta_{j}\sin(N_{j,k+1}\eta_{j}\cdot x))\|_{\infty}$ $\lesssim N_{j,k+1}^{n}.$
@@
$\|\nabla^{n}\mathbb{Q}F_{j,j^{\prime},k}\|_{\infty}$ $\lesssim N_{j\vee j^{\prime},k+1}^{n+1+\varepsilon}\ell_{k+1},$
@@
$\|\nabla^{n}\mathcal{E}_{k}^{3}\|_{\infty}$ $\lesssim\sum_{j,j^{\prime}}N_{j,k+1}N_{j^{\prime},k+1}N_{j\vee j^{\prime},k+1}^{-2}N_{j\vee j^{\prime},k+1}^{n+1+\varepsilon}\ell_{k+1}\lesssim\sum_{j}N_{j,k+1}^{n+1+\varepsilon}\ell_{k+1}.$
@@
$\begin{split}\|\nabla^{n}(\mathcal{J}_{k}^{1}+\mathcal{J}_{k}^{2}+\mathcal{E}_{k})\|_{\infty}&\leq\|\nabla^{n}\mathcal{J}_{k}^{1}\|_{\infty}+\|\nabla^{n}\mathcal{J}_{k}^{2}\|_{\infty}+\|\nabla^{n}\mathcal{E}_{k}^{1}\|_{\infty}+\|\nabla^{n}\mathcal{E}_{k}^{2}\|_{L^{\infty}}+\|\nabla^{n}\mathcal{E}_{k}^{3}\|_{\infty}\\ &\lesssim_{\varepsilon,m}\sum_{j}N_{j,k+1}^{n-1+\varepsilon}N_{J_{d},k}\big(e^{-\frac{1}{4}M_{j,k+1}^{2}t}+M_{j,k+1}^{-m}N_{J_{d},k}^{m}\big)\\ &\quad\,\,+\sum_{j^{\prime}<j}N^{n-1+\varepsilon}_{j,k+1}N_{j^{\prime},k+1}e^{-\frac{1}{4}N_{j,k+1}^{2}t}\\ &\quad\,\,+\sum_{j^{\prime}<j}N_{j,k+1}^{-1}N_{j^{\prime},k+1}N_{j,k+1}^{n-m}M_{j,k+1}^{m+\varepsilon}\\ &\quad\,\,+\sum_{j}\big(N_{j,k+1}^{n-1+\varepsilon}e^{-\frac{1}{4}N_{j,k+1}^{2}t}+N_{j,k+1}^{-2}M_{j,k+1}^{n+2+\varepsilon}\big)+\sum_{j}N_{j,k+1}^{n+1+\varepsilon}\ell_{k+1}.\end{split}$
@@
for $A$, $b$, and $m$ large enough. For $n=0$ the bound becomes worse on $0,t_{k}]$, as in Proposition 5.1. ∎
@@ 5.2. Bounds on the solution and residual
for some $0<\kappa<\alpha<1$. Moreover,
@@
$\lesssim t^{-1-\frac{n}{2}+\frac{1}{2}(1-b^{-1/J_{d}}-4\kappa)}.$
@@
$J_{2}=$ $\sum_{j}\Big(\sum_{k:N_{j,k}<t^{-1/2}}\int_{t^{\prime}}^{t}N_{j,k}s^{-\frac{1}{2}}ds+\sum_{t^{-1/2}\leq N_{j,k}\leq(t^{\prime})^{-1/2}}\int_{t^{\prime}}^{t}N_{j,k}s^{-\frac{1}{2}}e^{-N_{j,k}^{2}s}ds$
@@
$J_{2}\lesssim 1+\log_{A^{b-1}}((t^{\prime})^{-\frac{1}{2}}/t^{-\frac{1}{2}})$
@@ 6. Construction of the corrector
Let $U$ be any classical solution to the Navier–Stokes equations on the time interval $[0,T]$ which, once again, we replace with $T_{*},T-T_{*}]$ by translation. We may define
@@
The family of solutions $u^{(\sigma)}$ claimed in Theorem 1.7 will arise by modulating a frequency scale $N_{0}>0$ (see Section 7). Having constructed $v$ on $\mathbb{T}^{d}\times[0,\infty)$, we extend it by 0 to the full time interval $\mathbb{R}$, then rescale it to $v^{N_{0}}$, which we regard as a $2\pi/N_{0}$-periodic vector field on $\mathbb{R}^{d}\times[0,\infty)$. Here, for an arbitrary vector field on $\mathbb{R}^{d}$, we define the natural Navier–Stokes rescaling
@@
$\begin{split}\partial_{t}w-\Delta w+\mathbb{P}\operatorname{div}(w\otimes w+2(U^{1/N_{0}}+v)\odot w)&=-\mathbb{P}\operatorname{div}(f+2U^{1/N_{0}}\odot v),\\ w|_{t=0}&=0.\end{split}$
@@ 6.1. Semigroup estimates
Then we have the following product rule, which is immediate from the definitions and the estimate $\|gh\|_{C^{1,\kappa}}\lesssim\|g\|_{C^{1,\kappa}}\|h\|_{L^{\infty}}+\|g\|_{L^{\infty}}\|h\|_{C^{1,\kappa}}$.
@@
If $g,h\in X$, then $g\odot h\in Y$ with
$\|g\odot h\|_{Y}\lesssim\|g\|_{X}\|h\|_{X}.$
@@
$\|S(t,t^{\prime})\phi\|_{L^{\infty}(\mathbb{T}^{d})}+(t-t^{\prime})^{\frac{1}{2}}\|\nabla S(t,t^{\prime})\phi\|_{C^{\kappa}(\mathbb{T}^{d})}\lesssim_{\bar{C}}t^{-\frac{1}{2}}(t^{\prime})^{-1+\alpha}(t/t^{\prime})^{\epsilon}\|\phi\|_{Y}$
@@
$\begin{split}S(t,t^{\prime})\phi&=e^{(t-t^{\prime})\Delta}\mathbb{P}\operatorname{div}\phi(t^{\prime})-2\int_{t^{\prime}}^{t}e^{(t-s)\Delta}\mathbb{P}\operatorname{div}((U^{1/N_{0}}+v)(s)\odot S(s,t^{\prime})\phi)\,ds\\ &=e^{(t-t^{\prime})\Delta}\mathbb{P}\operatorname{div}\phi(t^{\prime})-2\int_{t^{\prime}}^{t^{\prime}\vee{\frac{t}{2}}}e^{(t-s)\Delta}\mathbb{P}\operatorname{div}((U^{1/N_{0}}+v)(s)\odot S(s,t^{\prime})\phi)\,ds\\ &\quad-2\int_{t^{\prime}\vee{\frac{t}{2}}}^{t}e^{(t-s)\Delta}\mathbb{P}\operatorname{div}((U^{1/N_{0}}+v)(s)\odot S(s,t^{\prime})\phi)\,ds\\ &=:I_{1}+I_{2}+I_{3}.\end{split}$
@@
Therefore we need an estimate for $\|v\|_{L^{1}([t^{\prime},t],t^{-\frac{1}{2}}dt;L^{\infty})}$ and $\|v\|_{L^{2}([t^{\prime},t];L^{\infty})}$. Using (5.19), we obtain
@@
$\begin{split}\nabla S(t,t^{\prime})\phi&=e^{(t-t^{\prime})\Delta}\nabla\mathbb{P}\operatorname{div}\phi(t^{\prime})-2\int_{t^{\prime}}^{t}e^{(t-s)\Delta}\nabla\mathbb{P}\operatorname{div}((U^{1/N_{0}}(s)+v(s))\odot S(s,t^{\prime})\phi)\,ds\\ &=e^{(t-t^{\prime})\Delta}\nabla\mathbb{P}\operatorname{div}\phi(t^{\prime})-2\int_{t^{\prime}}^{\frac{1}{2}(t+t^{\prime})}e^{(t-s)\Delta}\nabla\mathbb{P}\operatorname{div}((U^{1/N_{0}}(s)+v(s))\odot S(s,t^{\prime})\phi)\,ds\\ &\quad-2\int_{\frac{1}{2}(t+t^{\prime})}^{t}e^{(t-s)\Delta}\nabla\mathbb{P}\operatorname{div}((U^{1/N_{0}}(s)+v(s))\odot S(s,t^{\prime})\phi)\,ds\\ &=:I_{4}+I_{5}+I_{6}.\end{split}$
@@
$\|e^{t\Delta}\nabla^{m}\mathbb{P}g\|_{C^{r}}\lesssim t^{-\frac{m+r-s}{2}}\|g\|_{C^{s}}$ (6.4)
@@ 6.2. Fixed point argument
There exists $\varepsilon>0$ such that for all $\delta\in(0,\varepsilon)$ and $A\gg 1$, there exists $w\in B_{X}(0,\delta)$ with $u^{1/N_{0}}=U^{1/N_{0}}+v+w$ solving (1.1) on $0,\bar{T}]\times\mathbb{R}^{d}$. Further, choosing $\bar{C}$ and $N_{0}$ large depending on $U$ and $T$, the solution $u^{1/N_{0}}$ extends to a classical solution on $[0,T]$.
@@
$F(w)=-\int_{0}^{t}S(t,t^{\prime})(w\otimes w+f+2U^{1/N_{0}}\odot v)(t^{\prime})\,dt^{\prime}$
@@
$\|F(w)\|_{X}$ $\leq\sup_{t\in(0,\bar{T}]}\int_{0}^{t}\Big(t^{\frac{1}{2}(1-\alpha)}\|S(t,t^{\prime})(w\otimes w+f+2U^{1/N_{0}}\odot v)(t^{\prime})\|_{\infty}$
$\qquad+t^{1-\frac{\alpha}{2}}\|\nabla S(t,t^{\prime})(w\otimes w+f+2U^{1/N_{0}}\odot v)(t^{\prime})\|_{C^{\kappa}}\Big)\,dt^{\prime}$
$\lesssim_{\bar{C}}\sup_{t\in(0,\bar{T}]}\int_{0}^{t}\Big(t^{\frac{1}{2}(1-\alpha)}+t^{1-\frac{\alpha}{2}}(t-t^{\prime})^{-\frac{1}{2}}\Big)t^{-\frac{1}{2}}(t^{\prime})^{-1+\alpha}\left(\frac{t}{t^{\prime}}\right)^{\epsilon}$
$\qquad\qquad\cdot\|w\otimes w+f+2U^{1/N_{0}}\odot v\|_{Y}\,dt^{\prime}$
$\lesssim\|w\otimes w\|_{Y}+\|f\|_{Y}+\|U^{1/N_{0}}\odot v\|_{Y},$
choosing $\epsilon\in(0,\alpha)$. It follows from Lemma 6.1 that
$\|w\otimes w\|_{Y}\lesssim\|w\|_{X}^{2}.$
On the other hand, in view of the definition (6.2) and Proposition 5.3, we have
$\begin{split}\|f\|_{Y}&=\sup_{t\in(0,\bar{T}]}(t^{1-\alpha}\|f\|_{L^{\infty}}+t^{\frac{3}{2}-\alpha}\|\nabla f\|_{C^{\kappa}})\\ &\lesssim\epsilon_{0}\sup_{t\in(0,\bar{T}]}(t^{1-\alpha}(t^{-1+\alpha}+1)+t^{\frac{3}{2}-\alpha}(t^{-\frac{3}{2}+\alpha}+1))\\ &\lesssim\epsilon_{0},\end{split}$
and
$\begin{split}\|U^{1/N_{0}}\odot v\|_{Y}&=\sup_{t\in(0,\bar{T}]}(t^{1-\alpha}\|U^{1/N_{0}}\odot v\|_{L^{\infty}}+t^{\frac{3}{2}-\alpha}\|\nabla(U^{1/N_{0}}\odot v)\|_{C^{\kappa}})\\ &\lesssim 1/N_{0}\sup_{t\in(0,\bar{T}]}(t^{1-\alpha}\|v\|_{L^{\infty}}+t^{\frac{3}{2}-\alpha}\|v\|_{C^{\kappa}}+t^{\frac{3}{2}-\alpha}\|\nabla v\|_{C^{\kappa}})\\ &\lesssim 1/N_{0}\sup_{t\in(0,\bar{T}]}(t^{1-\alpha-\frac{1}{2}}+t^{\frac{3}{2}-\alpha-\frac{1}{2}(1+\kappa)}+t^{\frac{3}{2}-\alpha-\frac{1}{2}(2+\kappa)})\\ &\lesssim 1/N_{0}.\end{split}$
Therefore, for large enough $N_{0}>1$, we conclude
$\|F(w)\|_{X}$ $\leq C(\|w\|_{X}^{2}+\|f\|_{Y}+\|U^{1/N_{0}}\odot v\|_{Y})\leq C^{\prime}(\delta^{2}+\epsilon_{0}),$
@@
Taking $\delta>0$ small enough, we conclude the existence of a fixed point $w\in B_{X}(0,\delta)$ of $F$. This proves the claim about $u$ on $0,\bar{T}]$. For $N_{0}$ sufficiently large, it will be the case that $N_{0}^{2}(T-T_{*})>\bar{T}=\bar{C}$. To extend the lifespan, we argue that global solution theory can be applied starting from initial data $u(\bar{T})$. Observe that by (5.18) and the fact that $w\in X$, the perturbation $v(\bar{T})+w(\bar{T})$ can be made to be arbitrarily small by arranging $\bar{T}$ to be sufficiently large (which in turn is achieved with a sufficiently large choice of $N_{0}$ and $\bar{C}$). Moreover, $U^{1/N_{0}}(\bar{T})$ stays bounded in $BMO^{-1}$ by the scaling invariance. Taking $\bar{T}$ as large as needed depending on $U$, we conclude existence of $u$ up to time $T-T_{*}$ by (for instance) the stability theorem of Auscher, Dubois, and Tchamitchian [2]. ∎